OpenStudy (anonymous):

Help please. Where do I begin?... Lithium iodide has a lattice energy of −7.3×102kJ/mol and a heat of hydration of −793kJ/mol. Find the heat of solution for lithium iodide. Determine how much heat is evolved or absorbed when 10.0g of lithium iodide completely dissolves in water.

OpenStudy (anonymous):

First you have construct a cycle 

OpenStudy (anonymous):

therfore Heat of solution = heat of hydration - heat of lattice = -793 + 7.3 * 10^2 = E For one mole of Lithium iodid E amountof energy is release, then for 10 g how mch 1 = E 10/Molar mass of LiI = x E * 10/molarmass of LI I = x

OpenStudy (anonymous):

I think I got it!! Thank you! That diagram helped a lot. We didn't learn this in class so I'm having to try and figure out how to work it.